### v1

Abstract
The integration of Artificial Intelligence (AI) into Integrated Development Environments (IDEs) is reshaping software development, fundamentally altering how developers interact with their tools. This shift marks the emergence of Human-AI Experience in Integrated Development Environment (in-IDE HAX), a field that explores the evolving dynamics of Human-Computer Interaction in AI-assisted coding environments. Despite rapid adoption, research on in-IDE HAX remains fragmented which highlights the need for a unified overview of current practices, challenges, and opportunities. To provide a structured overview of existing research, we conduct a systematic literature review of 89 studies, summarizing current findings and outlining areas for further investigation.
Our findings reveal that AI-assisted coding enhances developer productivity but also introduces challenges, such as verification overhead, automation bias, and over-reliance, particularly among novice developers. Furthermore, concerns about code correctness, security, and maintainability highlight the urgent need for explainability, verification mechanisms, and adaptive user control. Although recent advances have driven the field forward, significant research gaps remain, including a lack of longitudinal studies, personalization strategies, and AI governance frameworks. This review provides a foundation for advancing in-IDE HAX research and offers guidance for responsibly integrating AI into software development.
Keywords:
Human-Computer Interaction Artificial Intelligence Integrated Development Environment Programming User Studies User Experience
MSC:
68N01 68T01 68U35
1 INTRODUCTION
The integration of AI into IDEs has gained significant traction in recent years, driven by the promise of enhanced developer productivity, improved software quality, and more efficient workflows (Ziegler et al., 2022; Izadi et al., 2024). Recent large-scale industry surveys show that 76% of respondents are either using or planning to use AI tools in their development process (Stack Overflow, 2024). Additionally, more than 50% of respondents note such benefits as reduced time spent searching for information, faster coding and development, quicker completion of repetitive tasks, and an overall increase in productivity (JetBrains, 2024).
Traditionally, Human-Computer Interaction (HCI) has focused on how users interact with software and tools. With AI now integrated into development workflows, this focus is shifting toward Human-AI Experience (HAX). Unlike conventional software, AI acts as an active collaborator rather than just a tool, changing how developers work and interact with their environment.
Given the growing reliance on AI-enhanced IDEs in software engineering (e.g., Coursor  [^1] Cursor - The AI Code Editor https://www.cursor.com/), it is crucial to merge and evaluate existing evidence in the field of in-IDE HAX. Doing so will help to identify recurring themes and gaps in the literature regarding the developer experience with AI assistance.
We previously conducted a literature review on in-IDE HAX (Sergeyuk et al., 2024), provided an initial exploration of the topic and highlighted its potential. However, it did not employ a systematic methodology and covered only 36 studies, resulting in less coverage of the rapid increase in both research and industry adoption of AI in IDEs. Currently, no comprehensive review merges this growing body of work.
This review aims to address this gap by systematically synthesizing recent research on in-IDE HAX, building on the foundation of our inital review. In this study, we adopt a broader scope and deeper analysis guided by Preferred Reporting Items for Systematic reviews and Meta-Analyses (PRISMA) framework (Moher et al., 2010).
Using our protocol, we identified 254 studies and thoroughly reviewed 89 that were relevant to in-IDE HAX. The goals of this review are to (a) merge the current state of knowledge regarding in-IDE HAX, (b) identify common themes and patterns in the literature, and (c) identify critical gaps that can guide future research efforts.
Our systematic literature review reveals key methodological trends, interaction paradigms, and the broader impact of AI integration in software development. Research in this domain predominantly employs qualitative (43%) and experimental (35%) approaches, often leveraging mixed methods to explore the evolving relationship between developers and AI. While the majority of studies focus on professional settings (82%), research on educational contexts (18%) remains limited, even though it is crucial for understanding long-term AI adoption and skill development.
Findings highlight two dominant AI interaction paradigms in IDEs: autocompletion -based assistance and conversational agents. Additionally, there is growing interest in hybrid models that combine real-time suggestions with interactive, context-aware support. Most studies ($\approx$53%) focus on GitHub Copilot, which primarily follows the autocompletion paradigm with chat as a supporting feature. However, other AI-powered development tools are also studied, some of which explore hybrid approaches. This demonstrates the variety of intelligent coding assistants.
The impact of AI on software development is complex. Although studies report productivity gains, they also highlight challenges such as verification overhead, automation bias, and over-reliance, particularly among novice developers. In AI-assisted code generation, concerns regarding correctness, security, and maintainability emphasize the need for explainability, verification mechanisms, and adaptive user control to ensure responsible adoption.
Our analysis highlights several areas that require further research. Many studies do not explicitly consider Software Development Lifecycle (SDLC) stages, limiting insights into AI ’s role beyond coding and debugging. Future work should examine AI assistance in earlier phases, such as requirements engineering and design, as well as later stages, such as testing and deployment. Furthermore, research on the long-term effects of AI-assisted development remains limited, particularly regarding skill retention and changes in collaboration. Personalization is another open challenge, with few studies exploring how AI can adapt to the expertise and workflows of individual developers. Finally, AI governance and ethical considerations, including bias mitigation and responsible adoption, require deeper investigation to ensure sustainable integration into software engineering.
Our study provides several contributions:
  • Comprehensive synthesis of existing research. We analyze 89 studies, categorizing them based on their research goals, methodological approaches, and key findings. This synthesis reveals dominant research themes, study contexts, and trends in HAX research.
  • Characterization of HAX research dimensions. We classify research efforts into three core aspects: (a) Design, focusing on how AI assistance is integrated into IDEs; (b) Impact, examining how AI affects developers’ workflows, productivity, and experience; and (c) Quality, assessing AI-generated code for security, comprehensibility, and adequacy.
  • Identification of future research opportunities. By analyzing the proposed future work in the studies, we identify under-explored topics, such as personalization of AI assistance, the long-term effects of AI adoption, and ethical considerations in AI-powered software development.
By addressing these aspects, our study provides a structured overview of in-IDE HAX research, informing both academic and industry efforts to design and evaluate AI-powered development tools.
2 METHOD
To answer our research questions, we set out to conduct a systematic literature review in accordance with the Preferred Reporting Items for Systematic Reviews and Meta-Analyses (PRISMA) framework (Moher et al., 2010).
The following Research Questions (RQs) guided our literature review:
  • RQ 1. What are the key findings in the field of Human-AI Experience within Integrated Development Environments? To answer RQ1, we systematically examined study contexts, goals, research questions, and key findings from published reports. That allowed us to determine key research themes in the field and define methodological trends.
  • RQ 2. What aspects of software development related to HAX in IDEs have been extensively studied, and which areas remain under-explored? To address RQ 2, we defined the stages of SDLC according to the Waterfall model (Alshamrani and Bahattab, 2015). The Waterfall model was chosen for its clear delineation of development stages, which facilitated systematic categorization. We also examined three key aspects of HAX identified by a previous literature review (Sergeyuk et al., 2024). Design of HAX draws on insights from interviews, literature reviews, experiments, and other domains to understand how HAX solutions are conceptualized and implemented. Impact explores whether and how HAX affects users, and investigates the nature of those effects. Quality focuses on the ability of AI to generate secure, comprehensible, and contextually appropriate code suggestions.
  • RQ 3. What are the potential directions for future research and development in the field of HAX in IDEs? To address RQ3, we analyzed the lists of Future Research directions proposed in the reviewed studies, identifying recurring themes and uncovering how studies build on each other’s findings.
2.1 Planning
Based on the best practices established in the field (Moher et al., 2010; Zhou et al., 2015; Carrera-Rivera et al., 2022; Kitchenham et al., 2010, 2009), we defined eligibility criteria to ensure a comprehensive and methodologically sound review. These criteria were guided by the research objectives and scope of this review. We report our protocol details below:
  • Timeframe: Studies published between 2022 and 2024 were included to ensure relevance to the advancements in AI-driven IDE features.
  • Language: Only studies published in English were included.
  • Publication Status: The review included peer-reviewed journal articles and conference proceedings. Recognizing the need to account for the rapid development in this field, preprints and unpublished dissertations were also considered if they provided unique insights directly relevant to the research objectives.
  • Population: This review considered studies in the field of computer science that focus on software developers or users engaging with AI-enhanced tools within IDEs.
  • Intervention: Studies were included if they investigated in-IDE HAX.
  • Comparison: Not applicable.
  • Outcomes: This review included studies that presented empirical findings, whether qualitative, quantitative, or mixed-methods.
  • Study Design: The review included original empirical studies.
During the systematic literature review’s planning stage, as recommended by the guidelines (Centre for Reviews and Dissemination (1995), UK), we established quality assessment criteria for studies that would be included after the initial screening phase. Our quality assessment questions were inspired by the recommendations of  Zhou et al. (2015):
  • Reporting: Is there a clear statement of the aims of the research?
  • Rigor: Is the study design clearly defined?
  • Credibility: Is there a clear statement of findings related to the aims of research?
  • Relevance: Is the study of value for research or practice?
Each question was accompanied by the following scale: 0 — No, and not considered; 0.5 — Partially; 1 — Yes. Following recommended guidelines, we set the cutoff for further inclusion in the review at a score higher than 2. These criteria ensured a systematic evaluation of study design, methodological soundness, and the validity of findings.
Furthermore, the data extraction form was defined during the planning stage. This form was guided by our research questions and the goal of assessing the representativeness of the included studies, given their varying venues and scopes. Therefore, in addition to the name of each report, the extraction form included the following fields; authors names and affiliation, DOI, publication date, venue, goal of the study, research questions, key findings, future work, scope (based on the taxonomy defined in previous work (Sergeyuk et al., 2024)), and SDLC stage.
To ensure coverage of a diverse range of scholarly sources, we selected several well-known digital libraries for our initial search. These included ACM Digital Library, DBLP, IEEE Digital Library, ISI Web of Science, ScienceDirect, Scopus, Springer Link, and arXiv. We included arXiv to capture insights from emerging fields, recognizing that valuable information can often be found in non-peer-reviewed papers. Additionally, given the prevalence of positive results in published literature, we included preprints and unpublished dissertations to reduce the risk of excluding studies with negative or null findings.
2.2 Identifying and Screening
In November 2024, after finishing the planning phase, we searched the aforementioned sources, using titles, keywords, and abstracts as the basis for our queries.
We used the following base string, adapting it for each source’s specific system requirements: (”AI Assistant” OR ”AI Companion” OR ”AI-Powered Programming Tool” OR ”Code Completion Tool” OR ”Coding Assistant” OR ”Copilot” OR ”Intelligent Code Assistant” OR ”LLM-Based Coding Assistant” OR ”LLM-Powered Coding Assistant” OR ”LLM4Code” OR ”Programming Assistant” OR ”Human-AI Experience” OR ”Human-AI Co-Creation” OR ”Human-AI Collaboration” OR ”Human-AI Collaboration” OR ”Human-AI Interaction”). The full list of search string variations with corresponding libraries can be found in Supplementary materials  [^2] The dataset can be accessed at https://doi.org/10.5281/zenodo.14976737.
The initial search overlooked 28 relevant papers known to us from our expertise in the field, so we added them manually. We also included 36 papers from our previous literature survey (Sergeyuk et al., 2024).
In total, 224 papers entered the screening phase of the systematic literature review. We identified and excluded 29 duplicates appearing across multiple databases. Next, we conducted an initial screening, applying the exclusion criteria and identifying 113 papers out of the study’s scope.
2.3 Assessing Eligibility
In the next stage of the systematic literature review, we assessed the eligibility of 82 full-text articles based on their quality. As described in Subsection 2.1, we applied four evaluation criteria: Reporting, Rigor, Credibility, and Relevance. Articles were accepted only if their combined score on these criteria exceeded 2.
The quality assessment was conducted independently by the second and third authors, followed by a discussion to reconcile any discrepancies. In cases of disagreement, the first author acted as an arbitrator. This structured process ensured consensus among the the reviewers for the final quality assessment. A total of 8 papers were excluded due to insufficient quality scores.
2.4 Backward Snowballing
To complement the systematic search, we performed backward snowballing (Wohlin, 2014) on papers deemed eligible for inclusion. First, we extracted citations from these papers and screened them against the report-related criteria outlined in Subsection 2.1. References that did not meet these criteria were excluded. We next removed any papers already identified in previous steps We assessed the remaining studies against the study-level inclusion criteria described in Subsection 2.1, excluding out-of-scope papers. After two rounds of this process, no new papers were identified, resulting in 18 new in-scope papers. These studies underwent the eligibility assessment process outlined in Subsection 2.3, yielding a final total of 89 papers to be included in the review.
2.5 Data Extraction
As mentioned in Subsection 2.1, from all 89 eligible papers, we extracted data about Authors, Authors’ Affiliation, DOI, Publication Date, Venue, Goal of the Study, Research Questions, Key Findings, Future Work, Scope and SDLC Stage. In the later stages of data analysis, we added a Context column to indicate whether HAX was studied in a professional or educational setting.
This work was conducted manually by the first author, with support from an AI-based system called Elicit  [^3] Elicit: The AI Research Assistant https://elicit.com/. This tool is designed to facilitate the extraction and summarization of scientific papers. It supported this work by formulating concise descriptions of each study’s goal, key findings, and future work. To ensure data integrity and accuracy, all tool-generated outputs were rigorously reviewed and cross-verified against the original papers. Information was included into the final review only after confirming its faithful representation of the original meanings. This hybrid approach ensured both efficiency and precision in the data extraction process.
Data extraction yielded a comprehensive table with 89 studies in it. We provide the full table in Supplementary materials (Sergeyuk et al., 2025), and present a shortened overview of the studies in Table LABEL:tab:overview.
Table 1: List of studies included in the review
Title Authors Date
“It would work for me too”: How Online Communities Shape Software Developers’ Trust in AI-Powered Code Generation Tools Cheng et al. (2024a) 2024
“It’s Weird That it Knows What I Want”: Usability and Interactions with Copilot for Novice Programmers Prather et al. (2023) 2023
A Case Study in Engineering a Conversational Programming Assistant’s Persona Ross et al. (2023b) 2023
A Case Study on Scaffolding Exploratory Data Analysis for AI Pair Programmers Zhou and Li (2023) 2023
A Large-Scale Survey on the Usability of AI Programming Assistants: Successes and Challenges Liang et al. (2024) 2023
A Mixed Reality Approach for Innovative Pair Programming Education with a Conversational AI Virtual Avatar Manfredi et al. (2023) 2023
A Study on Developer Behaviors for Validating and Repairing LLM-Generated Code Using Eye Tracking and IDE Actions Tang et al. (2024) 2024
A Transformer-Based Approach for Smart Invocation of Automatic Code Completion de Moor et al. (2024) 2024
A User-centered Security Evaluation of Copilot Asare et al. (2024) 2024
AI Tool Use and Adoption in Software Development by Individuals and Organizations: A Grounded Theory Study Li et al. (2024) 2024
AI-Powered Chatbots and the Transformation of Work: Findings from a Case Study in Software Development and Software Engineering Súße et al. (2023) 2023
An Analysis of the Costs and Benefits of Autocomplete in IDEs Jiang and Coblenz (2024) 2024
An Empirical Evaluation of GitHub Copilot’s Code Suggestions Nguyen and Nadi (2022) 2022
An Empirical Study of Code Search in Intelligent Coding Assistant: Perceptions, Expectations, and Directions Liu et al. (2024) 2024
An Empirical Study on Usage and Perceptions of LLMs in a Software Engineering Project Rasnayaka et al. (2024) 2024
An Exploratory Study on Upper-Level Computing Students’ Use of Large Language Models as Tools in a Semester-Long Project Tanay et al. (2024) 2024
Analyzing Prompt Influence on Automated Method Generation: An Empirical Study with Copilot Fagadau et al. (2024) 2024
Ansible Lightspeed: A Code Generation Service for IT Automation Sahoo et al. (2024) 2024
Anticipating User Needs: Insights from Design Fiction on Conversational Agents for Computational Thinking Penney et al. (2023) 2024
Are Prompt Engineering and TODO Comments Friends or Foes? An Evaluation on GitHub Copilot OBrien et al. (2024) 2024
Assessing the quality of GitHub copilot’s code generation Yetistiren et al. (2022) 2022
Better Together? An Evaluation of AI-Supported Code Translation Weisz et al. (2022) 2022
Can Developers Prompt? A Controlled Experiment for Code Documentation Generation Kruse et al. (2024) 2024
Case Study: Using AI-Assisted Code Generation In Mobile Teams Vasiliniuc and Groza (2023) 2023
CoLadder: Supporting Programmers with Hierarchical Code Generation in Multi-Level Abstraction Yen et al. (2023) 2023
Conversing with Copilot: Exploring Prompt Engineering for Solving CS1 Problems Using Natural Language Denny et al. (2023) 2023
Copilot Evaluation Harness: Evaluating LLM-Guided Software Programming Agarwal et al. (2024) 2024
Copilot for Xcode: Exploring AI-Assisted Programming by Prompting Cloud-based Large Language Models Tan et al. (2023) 2023
CoPrompt: Supporting Prompt Sharing and Referring in Collaborative Natural Language Programming Feng et al. (2023) 2024
Defendroid: Real-time Android code vulnerability detection via blockchain federated neural network with XAI Senanayake et al. (2024) 2024
Demystifying Practices, Challenges and Expected Features of Using GitHub Copilot Zhang et al. (2023a) 2023
Design Principles for Collaborative Generative AI Systems in Software Development Chen and Zacharias (2024) 2024
Designing PairBuddy—A Conversational Agent for Pair Programming Robe and Kuttal (2022) 2022
Developers’ Perspective on Today’s and Tomorrow’s Programming Tool Assistance: A Survey Kuang et al. (2024) 2024
Discovering the Syntax and Strategies of Natural Language Programming with Generative Language Models Jiang et al. (2022) 2022
Documentation Matters: Human-Centered AI System to Assist Data Science Code Documentation in Computational Notebooks Wang et al. (2022) 2022
Enhancing Code Completion with Implicit Feedback Jin et al. (2023) 2023
Evaluating Human-AI Partnership for LLM-based Code Migration Omidvar Tehrani and Anubhai (2024) 2024
Expectation vs. Experience: Evaluating the Usability of Code Generation Tools Powered by Large Language Models Vaithilingam et al. (2022) 2022
Exploring Interaction Patterns for Debugging: Enhancing Conversational Capabilities of AI-assistants Chopra et al. (2024) 2024
Exploring the Learnability of Program Synthesizers by Novice Programmers Jayagopal et al. (2022) 2022
Exploring the Problems, Their Causes and Solutions of AI Pair Programming: A Study on GitHub and Stack Overflow Zhou et al. (2025) 2024
From “Ban It Till We Understand It” to “Resistance is Futile”: How University Programming Instructors Plan to Adapt as More Students Use AI Code Generation and Explanation Tools such as ChatGPT and GitHub Copilot Lau and Guo (2023) 2023
Generation Probabilities Are Not Enough: Uncertainty Highlighting in AI Code Completions Vasconcelos et al. (2024) 2024
Github copilot in the classroom: learning to code with AI assistance Puryear and Sprint (2022) 2022
Grounded Copilot: How Programmers Interact with Code-Generating Models Barke et al. (2023) 2023
How Do Data Analysts Respond to AI Assistance? A Wizard-of-Oz Study Gu et al. (2023) 2024
How Novices Use LLM-Based Code Generators to Solve CS1 Coding Tasks in a Self-Paced Learning Environment Kazemitabaar et al. (2023) 2023
How Readable is Model-generated Code? Examining Readability and Visual Inspection of GitHub Copilot Al Madi (2022) 2022
IDA: Breaking Barriers in No-code UI Automation Through Large Language Models and Human-Centric Design Shlomov et al. (2024) 2024
Identifying the Factors That Influence Trust in AI Code Completion Brown et al. (2024) 2024
Improving Steering and Verification in AI-Assisted Data Analysis with Interactive Task Decomposition Kazemitabaar et al. (2024) 2024
In-IDE Generation-based Information Support with a Large Language Model Nam et al. (2023a) 2024
Investigating and Designing for Trust in AI-powered Code Generation Tools Wang et al. (2024) 2024
Investigating Explainability of Generative AI for Code through Scenario-based Design Sun et al. (2022) 2022
Investigating Interaction Modes and User Agency in Human-LLM Collaboration for Domain-Specific Data Analysis Guo et al. (2024) 2024
Is GitHub Copilot a Substitute for Human Pair-Programming? An Empirical Study Imai (2022) 2022
Is GitHub’s Copilot as Bad as Humans at Introducing Vulnerabilities in Code? Asare et al. (2023) 2023
Ivie: Lightweight Anchored Explanations of Just-Generated Code Yan et al. (2024) 2024
Lost at C: A User Study on the Security Implications of Large Language Model Code Assistants Sandoval et al. (2023) 2023
Methodology for Code Synthesis Evaluation of LLMs Presented by a Case Study of ChatGPT and Copilot Ságodi et al. (2024) 2024
Multi-line AI-assisted Code Authoring Dunay et al. (2024) 2024
Non-Expert Programmers in the Generative AI Future Feldman and Anderson (2024) 2024
On the Design of AI-powered Code Assistants for Notebooks McNutt et al. (2023) 2023
On the Robustness of Code Generation Techniques: An Empirical Study on GitHub Copilot Mastropaolo et al. (2023) 2023
Performance, Workload, Emotion, and Self-Efficacy of Novice Programmers Using AI Code Generation Gardella et al. (2024) 2024
Practices and Challenges of Using GitHub Copilot: An Empirical Study Zhang et al. (2023b) 2023
Practitioners’ expectations on automated code comment generation Hu et al. (2022) 2022
Practitioners’ Expectations on Code Completion Wang et al. (2023a) 2023
Productivity Assessment of Neural Code Completion Ziegler et al. (2022) 2022
Prompt Sapper: A LLM-Empowered Production Tool for Building AI Chains Cheng et al. (2024b) 2024
Reading Between the Lines: Modeling User Behavior and Costs in AI-Assisted Programming Mozannar et al. (2024a) 2024
Significant Productivity Gains through Programming with Large Language Models Weber et al. (2024) 2024
Slide4N: Creating Presentation Slides from Computational Notebooks with Human-AI Collaboration Wang et al. (2023b) 2023
Spellburst: A Node-based Interface for Exploratory Creative Coding with Natural Language Prompts Angert et al. (2023) 2023
Taking Flight with Copilot: Early insights and opportunities of AI-powered pair-programming tools Bird et al. (2022) 2023
The Impact of AI on Developer Productivity: Evidence from GitHub Copilot Peng et al. (2023) 2023
The Programmer’s Assistant: Conversational Interaction with a Large Language Model for Software Development Ross et al. (2023a) 2023
The RealHumanEval: Evaluating Large Language Models’ Abilities to Support Programmers Mozannar et al. (2024c) 2024
The Widening Gap: The Benefits and Harms of Generative AI for Novice Programmers Prather et al. (2024) 2024
Toward Effective AI Support for Developers: A survey of desires and concerns Khemka and Houck (2024) 2024
Towards Feature Engineering with Human and AI’s Knowledge: Understanding Data Science Practitioners’ Perceptions in Human&AI-Assisted Feature Engineering Design Zhu et al. (2024) 2024
Towards More Effective AI-Assisted Programming: A Systematic Design Exploration to Improve Visual Studio IntelliCode’s User Experience Vaithilingam et al. (2023) 2023
Transforming Software Development: Evaluating the Efficiency and Challenges of GitHub Copilot in Real-World Projects Pandey et al. (2024) 2024
Trust in Generative AI among students: An Exploratory Study Amoozadeh et al. (2024) 2024
Using AI Assistants in Software Development: A Qualitative Study on Security Practices and Concerns Klemmer et al. (2024) 2024
Using GitHub Copilot to Solve Simple Programming Problems Wermelinger (2023) 2023
Validating AI-Generated Code with Live Programming Ferdowsi et al. (2024) 2024
When to Show a Suggestion? Integrating Human Feedback in AI-Assisted Programming Mozannar et al. (2024b) 2024
3 RESULTS
Figure 1: Flow diagram of the study.
This section presents key findings from our review of the studies on human-AI interaction in software development. We analyze methodological trends, study contexts, AI tool design, AI’s impact on productivity and code quality, and emerging research directions. As described in Figure 1 and Section 2, we screened 254 papers, 100 of which were assessed for eligibility, with 89 included in the final review stage.
3.1 Methodologies
There are multiple methodological approaches used in human-AI interaction research. These methodologies can broadly be categorized into qualitative and quantitative approaches. Quantitative methods, in turn, can be further divided into survey-based studies and experimental research designs. Among 89 papers reviewed in this study, 12 employed a survey methodology, 31 followed an experimental design, and 38 utilized qualitative methods.
Qualitative studies included a range of approaches, from semi-structured interviews to more complex methods such as the Wizard of Oz paradigm (Robe and Kuttal, 2022), focus group interviews (Vaithilingam et al., 2023), and Grounded Theory-based analysis (Li et al., 2024). Some studies adopted a mixed-methods approach, combining multiple methodologies. For instance, an experimental design complemented by participant interviews (a format often referred to as a user study).
Additionally, 20 of the reviewed studies did not fit neatly into the qualitative-quantitative dichotomy. Examples include case studies in which the authors both describe their own experiences and provide quantitative metrics on system behavior (Ross et al., 2023b), quantitative analysis of system behavior using pre-defined datasets of typical user tasks (Nguyen and Nadi, 2022; Fagadau et al., 2024), and investigations of public discussions on platforms such as Reddit (Klemmer et al., 2024) and GitHub Issues (Zhou et al., 2025).
The sample sizes for different methodological approaches varied considerably. For interview-based studies, the sample size ranged from 7 to 61 participants (median = 16). Experimental designs involved sample sizes ranging from 17 to 214 participants, with the exception of two A/B studies, one with 535 participants (Mozannar et al., 2023) and another, conducted by Meta (Dunay et al., 2024), reporting “thousands” of participants. As expected, survey-based studies had the largest sample sizes, ranging from 68 to 2,047 participants (median = 507).
3.2 Context of studies
The studies analyzed in this review represent two primary contexts for working with AI in software development: professional (82%) and educational (18%). These contexts provide contrasting perspectives on the integration of AI tools, with professional studies focusing on real-world applications and educational studies exploring their role in learning environments.
The majority of professional studies emphasize the practical utility of AI tools. Most studies address coding (26%) and maintenance (9%), with limited exploration of requirements gathering (1%) or other SDLC stages. However, 45 studies (61%) do not specify the SDLC stage, indicating a gap in contextual precision.
Educational studies primarily focus on pedagogical insights. These studies often investigate how students use AI tools to learn programming or solve computational problems. While a few studies address coding and requirements gathering (15%), most (81%) do not specify the SDLC stage, aligning with the broader exploratory nature of educational research.
This distribution reveals notable gaps in the literature:
  • Lack of perspectives from non-adopters of AI tools for development.
  • Limited number of educational studies highlighting the need for a deeper exploration of AI ’s impact on CS education, especially as developers constantly learn and re-learn programming practices.
  • Lack of specification of the SDLC stage, suggesting an opportunity for future research to contextualize AI tool usage across all stages of the SDLC.
A subset of studies (5%) conceptualizes programming with AI as a form of pair programming, where AI acts as a collaborative partner. This perspective highlights the evolving view of AI as not just a tool but a collaborator, representing a growing trend in co-creative programming workflows and suggesting new opportunities and challenges for both practitioners and educators.
3.3 Design
One of the research areas that emerged from both our current review and our previous work (Sergeyuk et al., 2024) is the Design of HAX in IDE. Among the reviewed studies, approximately 31% investigate forms of AI integration into IDEs, with most of them (22 out of 28) explicitly examining how these forms of integration impact users.
The research presents two primary paradigms of AI-assisted interaction inside IDEs: autocompletion-based assistance and conversational agents. Autocompletion interfaces provide short, low-friction code completions, enabling developers to work seamlessly with inline suggestions. These approaches minimize cognitive load and support rapid decision-making without disrupting workflow (Vaithilingam et al., 2023; Yen et al., 2023). In contrast, conversational interactions — often embodied in chat-based assistants or virtual avatars — facilitate higher-level reasoning, iterative refinement, and complex problem-solving(Ross et al., 2023a; Robe and Kuttal, 2022; Gu et al., 2023). Rather than being opposing paradigms, these interaction modes are complementary (Weber et al., 2024), with studies advocating for hybrid models that seamlessly integrate autocompletion with interactive, context-sensitive conversations. In addition, some studies experiment with alternative modalities of interaction that go beyond these two dominant paradigms. For example, mixed-reality interfaces(Manfredi et al., 2023) introduce conversational AI in augmented pair programming environments, leveraging immersive learning experiences for skill development.
Due to the conversational nature of interaction with AI in IDE, prompting became a prominent object of investigations with 12% of studies focusing on this aspect. Across reviewed studies, prompt engineering is positioned as a strategic skill that shapes how developers and AI systems collaborate, affecting both individual productivity and team-based workflows. A central challenge in AI-assisted development is ensuring that AI-generated code aligns with user intent. Studies investigating prompt formulation (Fagadau et al., 2024; Denny et al., 2023) demonstrate that structured and well-contextualized prompts lead to more accurate and reliable output, while excessive detail can degrade performance. As AI tools become embedded in professional and educational workflows, collaborative prompt engineering is also gaining importance. Some works (Feng et al., 2024; Cheng et al., 2024b) introduce frameworks for shared prompt refinement, knowledge transfer, and workflow optimization, suggesting the need for modularity and re-usability in AI interactions.
Beyond interaction paradigms, around 10% of the studies explore design principles for AI-powered IDEs. These principles guide how AI should be embedded into developer workflows while maximizing usability, reliability, and efficiency:
  • Context Awareness: LLM-powered assistance significantly boosts task completion rates when it can effectively incorporate a developer’s code as context (Nam et al., 2023b).
  • Explainability & Transparency: Tools should provide context-aware explanations and highlight uncertain AI-generated code (Vasconcelos et al., 2024; Yan et al., 2024; Sun et al., 2022; Wang et al., 2024) to improve user trust.
  • User Control & Adaptability: Users need adaptive, fine-tuned to expertise or business context AI responses, and scaffolded AI guidance to prevent over-reliance (Prather et al., 2024; Jiang et al., 2022; Cheng et al., 2024b).
Among the specific implementations of AI for IDEs, GitHub Copilot is by far the most studied, with  53% of works investigating it. This concentrated attention on a single tool suggests a risk of overgeneralization, where findings may not fully translate to other AI-powered coding assistants, limiting broader design and implementation insights.
However, research also explores alternative AI-powered tools and approaches beyond Copilot, highlighting a growing landscape of intelligent development assistants. For example, IDA (Shlomov et al., 2024) introduces a no-code automation system tailored for non-programmers, while Prompt Sapper (Cheng et al., 2024b) supports AI chain development through modular and visual programming interfaces. Studies, such as CoLadder (Yen et al., 2023), investigate hierarchical prompt structures to refine interactions with code-generation models. PairBuddy (Robe and Kuttal, 2022) explores conversational AI as a simulated pair-programming partner. Domain-specific solutions, such as Slide4N (Wang et al., 2023b) for human-AI collaboration in computational notebooks and RealHumanEval (Mozannar et al., 2024c) for LLM performance evaluation in software engineering, illustrate the diversity of AI integrations in development workflows.
3.4 Impact
Another notable and prominent branch of research is dedicated to Impact of HAX in IDE with 82% of reviewed studies focusing on it.
The increasing integration of AI-powered tools in IDEs has reshaped the way developers engage with coding, debugging, documentation, and software design. AI ’s role extends beyond simple automation, influencing productivity, cognitive processes, and software engineering practices. Although AI has undoubtedly enhanced productivity in some contexts, its overall effects remain highly context-dependent and nuanced. Studies consistently highlight the dual nature of AI’s impact, providing both benefits and challenges.
Multiple studies have confirmed a notable boost in productivity when developers incorporate AI into their workflows. Potential benefits include a higher “ceiling” of productivity for experienced developers and faster onboarding for newcomers (Vasiliniuc and Groza, 2023). For instance, participants using GitHub Copilot completed their tasks up to 55.8% faster than control groups (Peng et al., 2023), while other investigators observed productivity gains ranging from 26% to 35% (Pandey et al., 2024). These improvements are partially attributed to fewer context switches (less time navigating external resources) and reduced boilerplate coding, enabling developers to offload repetitive tasks to AI while focusing on higher-level logic. However, gains are inconsistent: tasks involving proprietary or highly complex logic see only marginal benefits, as AI struggles with context-specific nuances.
Although AI-driven suggestions frequently accelerate development, they also require new verification efforts. When using AI tools, verifying suggestions, refining prompts, and reworking code generated by AI might take up to 50% of developers’ time (Mozannar et al., 2024a). This increased mental overhead arises because AI ’s output can be partially correct yet subtly flawed (Wermelinger, 2023), requiring careful review and, at times, stepwise interaction or re-prompting (Kazemitabaar et al., 2024). Such extra validation often mitigates the risk of blindly accepting erroneous suggestions — a phenomenon labeled “automation bias” (Al Madi, 2022).
Studies focusing on novice programmers reveal both benefits and pronounced risks (Prather et al., 2023; Puryear and Sprint, 2022). Beginners often solve tasks quicker with AI’s help. However, they can over-trust AI, losing sight of important foundational concepts (Prather et al., 2024). In certain cases, when system suggestions are misleading, novices drift away from correct solutions, as they lack the expertise to detect subtle errors (Zhou and Li, 2023). However, structured prompts and explanatory interfaces can help novices understand AI outputs better, fostering safer learning pathways (Kruse et al., 2024).
While most studies focus on short-term task performance, there is an emerging interest in longitudinal effects. Areas of concern include skill atrophy (if developers delegate too much to AI), shifts in collaboration norms among team members, and changes in programming practices.
3.4.1 Attitude
Some studies focus on the impact of HAX in the form of user attitudes toward AI systems integrated in IDEs.
Trust in AI-powered tools is heavily influenced by both the characteristics of AI outputs and user expertise (Amoozadeh et al., 2024; Brown et al., 2024). High-quality, contextually relevant suggestions improve trust, while issues such as inconsistent or overly complex recommendations hinder it. Other situational factors, such as task complexity or file type, also shape developer trust. Tools that fail to support developers in assessing the trustworthiness of outputs often lead to cautious or dismissive behaviors, while explainability layers, usage statistics, and control mechanisms have been proposed to enhance trust-building (Wang et al., 2024).
An important challenge related to trust in AI tools is over-reliance on AI-generated output. Over-reliance occurs when developers accept AI-generated outputs without sufficient validation, posing potential risks. Studies indicate that users tend to overestimate the reliability of AI-generated code (Al Madi, 2022). Conversely, some studies highlight under-reliance, where developers reject AI outputs due to complexity or lack of explainability, even when the code is correct (Sun et al., 2022). Live programming environments, which enable real-time validation, have shown potential in mitigating both over-reliance and under-reliance by lowering the cognitive effort needed for review (Ferdowsi et al., 2024).
3.5 Quality
Ensuring the Quality of AI-assisted code generation is a central concern of 21% of the papers in the reviewed literature. Across studies, quality is examined from multiple perspectives, including code correctness, security, and readability.
One recurring theme is the trade-off between efficiency and correctness. AI-assisted coding significantly accelerates development workflows, but at the cost of increased susceptibility to subtle errors, security vulnerabilities, and reduced maintainability (Sandoval et al., 2023; Tang et al., 2024; Pearce et al., 2021). Several studies suggest that while AI-generated code often appears syntactically valid, it can contain logical flaws, insecure patterns, or deviations from best practices (Pearce et al., 2021; Asare et al., 2024). As a result, effective quality control mechanisms, such as automated verification, static analysis, and model-aware debugging tools, are emphasized as critical to responsible adoption (Agarwal et al., 2024).
Beyond correctness, readability and maintainability emerge as key aspects of AI-generated code quality. Although AI can generate complex solutions rapidly, its output is not always optimized for human understanding. Studies highlight that AI-generated code may be less readable due to overly concise structures, unconventional variable naming, or lack of meaningful comments (Al Madi, 2022). This raises concerns about long-term maintainability, particularly in team-based software development, where code legibility is as important as functional correctness. Research suggests that integrating AI-powered refactoring and explanation tools can help mitigate these challenges, enhancing both short-term usability and long-term code quality (Yan et al., 2024).
Security poses an additional concern: up to 36% of vulnerabilities in AI-assisted code originate from the models themselves, pointing out the risks of replicating insecure patterns from training data (Sandoval et al., 2023). However, programmers remain a major source of vulnerabilities, emphasizing the need for cautious oversight of AI suggestions.
Finally, quality in AI-assisted development is not just a technical issue but also a human and organizational challenge. Research highlights that quality outcomes are influenced by how developers interact with AI tools: whether they verify AI-generated code, how they prompt AI models, and whether they develop strategies for AI-assisted debugging (Weber et al., 2024; Mozannar et al., 2024c; Sandoval et al., 2023; Al Madi, 2022; de Moor et al., 2024). Best practices for improving AI-assisted development quality extend beyond tool design and include education, training, and workflow adaptation.
While LLM-powered tools improve software development, maintaining consistently high-quality code remains a challenge. Ensuring reliability requires better alignment with developers’ needs and robust verification mechanisms (Tang et al., 2024; Agarwal et al., 2024). Studies highlight the need for strategies such as filtering suboptimal suggestions, improving contextual prompting through docstrings and function names, and leveraging adaptive personalization techniques (Mozannar et al., 2024c; de Moor et al., 2024).
Future research calls for more diverse benchmarks, longitudinal studies assessing real-world use, and broader quality evaluations that account for maintainability, reliability, and security (Sandoval et al., 2023; Pearce et al., 2021). Advancing AI-assisted programming will depend on integrating these refinements to ensure not only efficiency, but also long-term software quality.
3.6 Future work proposed by the studies
By analyzing the future work proposed by each study alongside the goals of others, we identified alignment patterns and gaps that highlight both the progress made and the opportunities missed.
Some future work suggestions align with the goals of other studies, but these connections are not always explicit, reflecting a fragmented research landscape. Common themes like productivity, security, and usability appear across studies but are addressed in isolated contexts without direct cross-referencing. This lack of continuity limits the field’s ability to build cumulative knowledge effectively.
A significant portion of the suggested directions for future research remains unaddressed, indicating how young the field is and highlighting a variety of opportunities for the academic community. Key gaps persist, particularly in studying personalization opportunities, adoption barriers, and cross-context comparisons. Moreover, many studies suggest conducting longitudinal research or involving larger sample sizes. Furthermore, ethical challenges such as bias mitigation, misuse prevention, and AI governance in software development remain largely unresolved.
Addressing these gaps through deliberate cumulative and interdisciplinary research will not only advance the field but also ensure its continued relevance in shaping AI-enhanced software development.
4 DISCUSSION
This systematic review revealed a growing body of research on AI-assisted development that intersects three key themes: Design, Impact, and Quality. From the Design perspective, the most important areas of research are related to transparency and simplicity of the developing AI tools. Design research focuses primarily on autocompletion and conversational agents, with growing interest in hybrid models. From the Impact perspective, AI improves productivity, especially for repetitive tasks, but also increases verification efforts and risks automation bias. From the Quality perspective, AI-generated code is found to be efficient but prone to errors, security risks, and maintainability issues.
The reviewed studies used qualitative, quantitative, and mixed methods designs, with a considerable part of the research falling outside the qualitative-quantitative division (e.g.,, development of the new AI tools). It should also be noted that 82% of the studies focus on using AI tools in professional settings, while 18% explore AI in educational contexts.
Many reviewed studies focus on how AI affects coding tasks (e.g.,, autocompletion for boilerplate code, conversational agents for complex problem solving) rather than why these tools succeed or fail in particular contexts. Aligning the two perspectives can clarify both the Design of the IDE integrations and the Impact on the practitioners’ everyday workflows (Weber et al., 2024; Robe and Kuttal, 2022). Rather than treating AI as a mere feature, researchers can adopt HCI-informed frameworks that consider user trust, mental models, and organizational culture as central to tool adoption.
It is evident that AI-assisted tools can boost code production speed (Peng et al., 2023; Pandey et al., 2024), but the depth of these gains appears to be strongly context-dependent. Many studies document the promise of conversational agents and autocompletion (Vaithilingam et al., 2023; Ross et al., 2023a), but the success of these paradigms depends on matching the right interaction mode with the right development stage.
The HCI theories on mixed-initiative design (Amershi et al., 2014) recommend interfaces that dynamically adjust the level of AI control, factoring in the user’s expertise and the complexity of the tasks. Novices, for example, may improve their immediate performance yet fail to internalize key programming concepts (Zhou and Li, 2023). Studies of educational contexts (Puryear and Sprint, 2022) suggest that real gains occur when learners receive structured prompting, consistent feedback, and explicit guidance in validating AI outputs (Prather et al., 2024; Kruse et al., 2024). Moreover, advanced developers, while benefiting from speed-ups, must spend a substantial portion of time auditing and refining AI outputs to ensure correctness in highly domain-specific tasks (Mozannar et al., 2024a).
Prior research highlights the risk of subtle flaws, security vulnerabilities, and poor maintainability (Sandoval et al., 2023; Pearce et al., 2021). Despite the existence of automated checks and static analysis, most quality control remains a human responsibility (Mozannar et al., 2024a; Tang et al., 2024). Therefore, inline suggestions, while streamlining coding, warrant the need for transparent AI behavior and meaningful explanations to mitigate “automation bias” (Al Madi, 2022) and acceptance of incorrect code. Effective trust calibration can involve real-time feedback loops (e.g.,, automated test suites) or explicit disclaimers about the AI ’s uncertainty (Kocielnik et al., 2019). In line with HCI work on explainable AI, embedding contextual clues (e.g.,, highlighting AI confidence levels or providing direct links to relevant training data) may calibrate trust more effectively (Kocielnik et al., 2019). Such design elements should also encourage user control, letting developers accept or reject suggestions with a clear sense of the trade-offs (Prather et al., 2024).
The ethical dimensions of AI in software development remain underexamined. For instance, non-native English speakers or beginners may struggle with complex prompts, inadvertently reinforcing existing inequities in open-source communities or corporate teams. Adopting AI within professional software teams shifts norms around collaboration, with some studies likening AI tools to “pair programmers” (Robe and Kuttal, 2022). This conceptual shift raises questions about division of labor and collective ownership of code: developers can offload routine checks to the AI, but may lose deeper engagement with design patterns or system architecture. It also introduces new roles, such as “prompt engineer” or “model tutor”, highlighting the need for training team members not only to code, but also to manage and interpret model outputs (Feng et al., 2024; Cheng et al., 2024b). Integrating AI tools at scale requires policies that balance innovation with accountability. In fields like healthcare or finance, AI-driven coding missteps could carry elevated risk (Pearce et al., 2021). Drawing on sociotechnical systems perspectives in HCI, organizations may need to formalize acceptance criteria for AI-generated suggestions, define clear roles for oversight, and establish audit trails for code provenance (Amershi et al., 2019).
Our findings align with a broader meta-analysis (Vaccaro et al., 2024) of 106 studies on human–AI collaboration, which reveals nuanced performance trade-offs depending on a task type, human expertise, and the relative strengths of humans and AI. The meta-analysis found that human–AI combinations often perform worse than the best individual performer (Hedges’ $g=-0.23$), with performance losses in decision-making tasks and gains in creative content creation. Although these results do not yet seem to be the same for coding tasks in real-world applications, both analyses emphasize the role of user expertise in mediating outcomes. Novices struggle with interpreting AI outputs, often leading to reliance on incorrect suggestions, while experienced developers leverage AI more effectively for repetitive or well-defined tasks. Importantly, both highlight the heterogeneity of impacts, with collaboration being highly task-specific and sensitive to how AI tools align with human workflows. To improve human–AI systems, the meta-analysis suggests balancing human and AI capabilities, while our review emphasizes domain-specific considerations: enhancing explainability, tailoring AI tools to user expertise, and refining prompting mechanisms to reduce errors. These findings jointly point out the importance of designing adaptive, user-centered systems that align with task contexts and user needs to maximize the complementary strengths of humans and AI.
Our systematic review of the literature indicates one recurring issue in the experimental part of the research employed within the HAX field. We observed that studies often feature small sample sizes (median being 17 participants) without justification for the number of participants, and there is a total lack of preregistered research. Given the rapid pace of progress in the area, it is also challenging to assess the potential replicability of ongoing research and to address problems such as publication bias. The replicability crisis has become the most discussed regarding behavioral and social sciences. However, concerns regarding research rigor are also prevalent within Computer Science, particularly in areas such as AI and Machine Learning (Henderson et al., 2019). To address these challenges, we believe that the HAX research community should consider adopting practices from neighboring fields. Particularly psychology, which has faced a replicability crisis but has made significant strides toward overcoming it (Nosek et al., 2022). We propose the widespread adoption of practices such as preregistration, the routine use of statistical power analysis, and the implementation of blinding and randomization procedures where possible. Additionally, embedding transparency checklists as standard practice can guide researchers in documenting their methodology, statistical tests, data exclusion criteria, and data-sharing protocols. By integrating these practices into the peer-review process and rewarding transparency through incentives, the HAX community can align with other fields that have successfully improved replicability and credibility.
4.1 Threats to the Validity
While gathered findings provide valuable insights, in any scientific study, it is essential to consider potential threats to validity, which are factors that can affect the accuracy and generalizability of the research findings.
Sampling Bias: Despite rigorous efforts to include well-known libraries and refine the search string, the possibility of sampling bias remains. Although our approach aimed to minimize bias, the inherent challenge of capturing every relevant work persists. That is why we provide our search protocol to mitigate this threat.
Temporal Bias: The chosen time frame (papers published between 2022 and 2024) introduces a potential temporal bias, excluding earlier works. This decision was driven by the intention to focus on contemporary developments following the advent of LLMs. While acknowledging potential temporal limitations, we aim to capture the latest advancements and trends in this rapidly evolving field.
Source Reliability: Inclusion of non-peer-reviewed papers from ArXiv introduces concerns regarding the reliability of findings, given the absence of a formal peer-review process. Recognizing this limitation, we deemed it necessary to consider insights from ArXiv due to the dynamic and rapidly evolving nature of the field. Thoroughly examining titles, abstracts, and full texts ensured that all included works contributed meaningfully to our survey. Moreover, the open publication environment of ArXiv encourages the publication of negative or null results, contributing to a more balanced representation of the research outcomes.
Interpretation Bias: Analyzing a large amount of information can introduce interpretation bias and impact the way studies are categorized. While acknowledging this complexity, we emphasize transparency and provide the entire dataset (Sergeyuk et al., 2025) for the readers.
Challenges of industry-related research: It is important to acknowledge that, overall, HAX research faces unique challenges due to its close ties with the industry. One of the most significant challenges of conducting empirical studies in industrial settings is the ”time factor.” Industrial sponsors and participants often have short time horizons. They expect valuable feedback within 1-2 months of research, so that they could make timely business decisions. Although industrial sponsors and project managers are interested in rigorous results, practical constraints sometimes make it challenging to prioritize statistical significance. Analyzing immediate trends and patterns can be sufficient for them to assess the performance of their projects, teams, or technologies. By fostering a collaborative environment, we can develop strategies that balance the demands of industrial collaboration with the standards of academic research.
5 CONCLUSION
Overall, our findings show that AI-assisted development has already begun transforming the nature of software engineering work, but the journey is far from complete. The core challenges revolve around (1) aligning AI capabilities with human expertise, (2) embedding transparency and accountability into developer tools, and (3) cultivating robust validation practices. From an HCI point of view, these challenges emphasize the importance of user-centered design, inclusive experimentation, and empirically-grounded evaluations that extend beyond short-term productivity metrics. We believe that sustained collaboration among researchers, practitioners, and end users will be essential for guiding innovation, refining best practices, and ultimately unlocking the full potential of AI-assisted development.
Acknowledgements.
This work was conducted as part of the AI for Software Engineering (AI4SE) collaboration between JetBrains and Delft University of Technology. The authors gratefully acknowledge the financial support provided by JetBrains, which made this research possible.
Conflict of interest
The authors have no conflict of interest with the content of this article.
Data availability
All data are available in our supplementary materials Sergeyuk et al. (2025).